OpenStudy (anonymous):

Write the following as a single square root: the square root of 8 times the square root of 3. No decimal answers, please!

OpenStudy (anonymous):

sqrt8 * sqrt3 = sqrt24

jimthompson5910 (jim_thompson5910):

you can simplify that to get \[\sqrt{24}=\sqrt{4*6}=\sqrt{4}\sqrt{6}=2\sqrt{6}\] So \[\sqrt{24}=2\sqrt{6}\] which means that \[\sqrt{8}\times\sqrt{3}=2\sqrt{6}\]

OpenStudy (anonymous):

what about the square root of 15 divided by the square root of 3? would it just be the square root of 5...?

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks! :)